determine the axis of symmetry of the quadratic function f(x) = 3x^2 - 5x - 1
Answer by nerdybill(7384)

f(x) = 3x^2 - 5x - 1
axis of symmetry:
x = -b/(2a)
x = -(-5)/(2*3)
x = 5/(2*3)
x = 5/6